How many grams of HCl are produced when 15.2 mol H2O reacts?

HCl produced: g


How many moles of CO2 are produced when 76.5 g HCl is produced?

CO2 produced: mol


How many grams of C12H4Cl6 are reacted when 100.25 g CO2 is produced?

C12H4Cl6 reacted: g


How many grams of HCl are produced when 2.5 kg C12H4Cl6 reacts?

HCl produced:g


1
Expert's answer
2021-11-16T23:52:02-0500

a)

Moles of H2O = 10.0 moles of O2 x 2 moles of H2O / 23 moles of O2 = 0.870 moles

b)

Moles of HCl = 15.2 moles of H2O x 12 moles of HCl / 2 moles of H2O = 91.2 moles

Mass of HCl = 91.2 moles x 36.46 g / mol = 3.33 x 10^3 g

c)

Moles of HCl = 76.5 g x 1 mol / 36.46 g = 2.10 moles

Moles of CO2 = 2.10 moles of HCl x 24 moles of CO2 / 12 moles of HCl = 4.2 moles

d)

Moles of CO2 = 100.25 g x 1 mol / 44.01 g = 2.28 moles

Moles of C12H4Cl6 = 2.28 moles of CO2 x 2 moles of C12H4Cl6 / 24 moles of CO2 = 0.19 moles

Mass of C12H4Cl6 = 0.19 moles x 360.88 g / mol = 68.6 g